# A Review of Barren Plateaus in Variational Quantum Computing

### v1

@@ III.3 Methods for analyzing variance
At the simplest level, one can heuristically determine if the loss or its gradients are concentrating by evaluating them at randomly sampled parameters and computing how the statistical variance scales with the number of qubits [8, 26, 18]. Similarly, one can also compute other quantities which have been proposed as surrogates and thus are able to study if a BP occurs. These include the landscape’s information content [27], the local state’s entanglement [28], or the landscape’s Fourier coefficients [29, 30, 31]. While such approaches can provide a rough estimate of the scaling, care must be taken as they suffer from statistical errors (as we measure the loss a finite number of times), and are generally unable to distinguish between the two types of BPs.
A preferred, but more theoretically demanding, method for detecting BPs is by analytically studying the scaling of the variance. Here, one needs to make assumptions regarding the PQC to make the problem more tractable. For instance, if $\mathcal{U}_{\bm{\theta}}$ is a deep unitary circuit whose distribution of unitaries forms at least a $2$-design [32] over a Lie group (see Boxes III.3-IV.2), one can use representation theory to exactly evaluate the variances [8, 33, 34, 35, 36, 18, 37, 38, 39, 40, 41, 42, 43] via Weingarten calculus [44, 45]. Instead, for shallow circuits that do not form designs, one can instead assume that the local gates are sampled from a local group, thus mapping the variance evaluation to a Markov chain-like process which can be analyzed analytically [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], numerically via Monte Carlo [57] and tensor networks [58, 59], or studied via XZ-calculus [60]. One can also take an iterative approach to explicitly integrate over the random parameters within a given angle range [61, 62, 54, 24] or exploit tools from Floquet theory [63, 64]. Then, if $\mathcal{U}_{\bm{\theta}}$ is a noisy, one can study the variance by making assumptions regarding to the type of noise acting throughout the PQC [25, 65, 66, 67, 68].
@@ IV Origins of BPs
While theoretical and numerical calculations can be used to determine the flatness of the landscape, these only identify a symptom of the variational quantum computing scheme. In this section we discuss recent findings which argue that BPs appear as a form of curse of dimensionality, as well as showcase how this phenomenon reveals itself under different scenarios. By understanding the underlying connection between all of them, we will see that one should be able to predict if a model will exhibit an exponentially concentrated loss.
@@ IV.1 A curse of dimensionality
Now, of course, this does not mean that comparing any pair of exponentially large objects will always lead to exponentially small signals. Indeed, conventional quantum algorithms [74] manipulate vectors in the exponentially large Hilbert space by carefully choreographing the patterns of constructive and destructive interference of the probability amplitudes [75]. Indeed such algorithms have even been adapted to establish advantage proofs for quantum machine learning [76, 77]. Yet, there are not many such algorithms in the variational setting as their construction requires carefully and purposely designed circuits with appropriate inductive biases for the task they are trying to solve. How well variational quantum algorithms can be coerced into finding and exploiting such structure is very much an open question.
@@ IV.2 Circuit expressiveness
Perhaps the main culprit responsible for the curse of dimensionality is the PQC as it has been shown that its expressive power (see Box IV.2) is directly linked to BPs. Roughly speaking, the larger the space the PQC explores in an unbiased manner, the more it is prone—when randomly initialized—to lead to input states that have exponentially small inner products with the measurement operator. At a first glance, this result seems counter-intuitive as we would ideally like to choose a PQC that is as expressive as possible to try to guarantee that a solution for Eq. (4) exists within the parameter landscape. However, the fact that PQCs with unbiased highly expressive features lead to BPs should be understood as a form of No-Free-Lunch theorem: The more uniformly expressive the model (i.e., the less inductive biases), the harder to train. Said otherwise, there is no one PQC to rule them all.
@@
[t] Box 6: Expressive power of the PQC The expressive power of a unitary PQC quantifies the breath of unitaries that the circuit is likely to produce in an unbiased manner. For instance, one can quantify the potential, or ultimate expressiveness of the PQC by its dynamical Lie algebra $\mathfrak{g}$ defined in Box III.3—since any unitary $U(\bm{\theta})$ produced by the PQC belongs in the dynamical Lie group $G=e^{\mathfrak{g}}$. Here, the larger the dimension of the algebra, the more expressive the circuit is, with controllable circuits being defined as those for which $\mathfrak{g}=\mathfrak{su}(2^{n})$. Of course, shallow circuits will not be able to cover the whole dynamical Lie group $G$, as they can only produce unitaries in some subset $\mathcal{S}\subseteq G$. To deal with this situation, we can instead consider the current expressiveness which is quantified by the difference between $\tau_{\mathcal{S}}^{(t)}$ and $\tau_{G}^{(t)}$. When those operators match, we say that the PQC forms a $t$-design [32]. Importantly, if $|\tau_{\mathcal{S}}^{(2)}-\tau_{G}^{(2)}|\leqslant\varepsilon$ (in operator norm) then [37] $|{\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]-{\rm Var}_{G}[\ell_{\bm{\theta}}(\rho,O)]|\in\mathcal{O}(\varepsilon\lVert O\rVert_{1})\,,$ (12) showing that we can quantify how much the variance of a shallow circuit deviates from that of a Haar random PQC over $G$ via $\varepsilon$.
The previous connection between expressiveness and BPs has been mathematically formalized for unitary [8, 46, 78, 18, 37, 38, 39, 79] and noisy [80] circuits, as one can show that the more expressive a circuit is, the more the loss can concentrate. A similar statement can be made in terms of the PQC’s capacity [81, 82], a measure of how many functions it can represent, as the larger the model’s capacity is, the more its loss can concentrate [83].
@@
In particular, let us review the recent approach in Refs. [37, 38, 39] for obtaining an exact analytical form for the loss function variance of deep unitary PQCs. By deep we refer to the regime where the number of layers is large enough to guarantee that the distribution of unitaries corresponding to random parameter choices forms a $2$-design [32]. That is, that it approximately matches the Haar distribution on $e^{\mathfrak{g}}$ (where $\mathfrak{g}$ is the circuit’s dynamical Lie algebra (DLA), defined in Box III.3) up to second moments with small additive approximation error. Assume for simplicity that either $\rho$ or $O$ belong to a group-module $\mathcal{M}\subseteq\mathcal{B}(\mathcal{H})$ that is non-trivial under the adjoint action of the circuit. (We recall that a group-module is a vector space that holds a representation for the action of a group. In the case of $\mathcal{B}$, $\mathcal{M}$ is a subspace that is closed under the adjoint action of $e^{\mathfrak{g}}$, i.e., if $A\in\mathcal{M}$, then $UAU^{\dagger}$ also belongs to $\mathcal{M}$ for any $U\in e^{\mathfrak{g}}$  [^1] Technically speaking, group modules are irreducible representations of the dynamical Lie group over the vector space $\mathcal{B}(\mathcal{H}_{0})$). One finds
@@ IV.3 Input states and measurements
While highly expressive PQC bear significant fault in leading to BPs, there exist cases where circuits with small expressiveness can still exhibit BPs depending on who the initial state and measurement operator are. For instance, let us consider Eq. (13) again, and assume that the ensuing DLA is not exponentially large, i.e., that we have $\dim(\mathfrak{g})\in\mathcal{O}(\operatorname{poly}(n))$. This case will arise when $\mathcal{U}_{\bm{\theta}}$ has strong inductive biases such as encoding some symmetries (see below for specific examples). Now, despite using a circuit with a non-exponentially large underlying algebra, the loss can still be exponentially concentrated. For instance, if we pick $O$ belonging to an exponentially large module (i.e., small algebras can admit exponentially large modules), then one will find exponential concentration from the denominator in Eq. (13). However, if $O$ belongs to a module whose dimension grows only polynomially with $n$, such as $iO\in\mathfrak{g}$, we will have that the variance’ denominator $\dim(\mathcal{M})=\dim(\mathfrak{g})$ only vanishes polynomially with $n$, meaning that the circuit expressiveness cannot be responsible for a BP. In fact, in this case the loss is obtained by comparing objects that live in a polynomially large subspace, thus potentially avoiding the curse of dimensionality. However, a deterministic BP is still possible if the norm of the projection of $\rho$ into the module is exponentially small, i.e., if $\mathcal{P}_{\mathcal{M}}(\rho)\in\mathcal{O}(1/2^{n})$. In Box IV.3 we present an example of an inexpressive circuits for which a BP can arise either because $O$ belongs to an exponentially large $\mathcal{M}$, or because $O$ is taken from a polynomial module but the state is ill-aligned.
@@
[t] Box 7: Single-qubit rotation PQCs can have BPs Consider a unitary circuit composed of general single qubit gates as $U(\bm{\theta})=\prod_{j=1}^{n}U_{j}(\bm{\theta}_{j})$, where $U_{j}(\bm{\theta}_{j})$ acts only on the $j$-th qubit. Clearly, this circuit is fairly inexpressive as it doesn’t even generate entanglement. Moreover, its DLA is simply given by $\mathfrak{g}=\oplus_{j=1}^{n}\mathfrak{su}(2)$, and thus of dimension $\dim(\mathfrak{g})=3n\in\mathcal{O}(\operatorname{poly}(n))$. Yet, we can see that if $O=Z^{\otimes n}$, then the operator belongs to an exponentially large module composed of all $3^{n}$ Pauli operators acting non trivially on all the qubits. Assuming that each $U_{j}(\bm{\theta}_{j})$ forms a $2$-design over $\mathbb{SU}(2)$ [37], a direct calculation reveals that for all $\rho$, ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]\in\mathcal{O}(1/2^{n})$. However, if $O=Z_{1}$, then the measurement belongs to a module composed of the $3$ Pauli operators acting on the first qubit. Now, Eq. (13) leads to ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]=2({\rm Tr}[\rho_{1}^{2}]-1/2)/3$, with $\rho_{1}$ the reduced state of $\rho$ on the first qubit. The choice of measurement operator can lead to a BP. When $O=Z^{\otimes n}$ the loss is obtained by comparing operators that can move in exponentially large modules, while for $O=Z_{1}$, the circuit can only move in the subspace of operators acting only on the first qubit. However, while the PQC and measurements are not responsible for BPs, the initial state can lead to an exponentially concentrated loss. In particular, if the state is too entangled and $({\rm Tr}[\rho_{1}^{2}]-1/2)\in\mathcal{O}(1/2^{n})$ (which occurs if $\rho$ satisfies a volume law of entanglement [102]), then a deterministic BP occurs.
@@ V Barren-plateau-prone architectures
In the previous section we have provided some intuition regarding the origins of BPs. In this section we review some BP-prone architectures. We note that in most cases it has been shown that deep versions of these architectures can lead to exponentially concentrated losses, as their DLAs are exponentially large. However, this does not mean that they cannot be useful in their shallow forms, or when leveraging smart initialization strategies.
@@ V.1 Hardware efficient ansatz and deep unstructured circuits
Hardware efficient ansatz [108] is a generic term employed for unstructured PQCs composed of single-qubit rotation gates interleaved with (fixed or parametrized) entangling gates. The term hardware-efficient spans from the fact that the gates of the circuit are chosen not because they encode some strong inductive bias from the problem one is trying to solve, but because they are easy to implement on the device (e.g., one only picks entangling gates that follow the device connectivity to avoid compiling overhead [87, 109, 110]). Under quite general conditions, hardware efficient ansatzes will be universal  [18], meaning that their DLA is $\mathfrak{g}=\mathfrak{su}(2^{n})$. As such, deep versions of these circuits are as expressive as possible, and therefore will exhibit BPs irrespective of the initial state or measurement operator [111, 33, 8, 46, 112, 113, 114]. Here it is important to note that several standard architectures exist in the literature, such as entangling gates acting on a brick-like fashion on alternating pairs of qubits or on a ladder. While these can indeed exhibit different performances in their shallower forms, they will ultimately lead to a BP if the number of layers is too large.
@@ V.2 Some problem-inspired ansatzes
Just like hardware efficient is used for a wide-range of problem-agnostic circuits, problem-inspired ansatzes is a blanket term used for models which embed inductive biases about the problem at hand into the circuit architecture itself. In most cases, problem-inspired ansatzes change the circuit expressiveness by using particular sets of gates, entangling gate topologies, or by correlating the parameters to respect some symmetry of the task. Here, one can generally expect that $\mathfrak{g}$ will be smaller than $\mathfrak{su}(2^{n})$ [37]. While we will present below circuits which can in fact sufficiently reduce the circuit expressive power to potentially avoid BPs, we here discuss problem-inspired architectures that can still have exponentially concentrated loss functions despite their inductive biases.
@@
Quantum Alternating Operator Ansatz (QAOA). The QAOA [117, 118] is widely used to variationally attempt to solve combinatorial optimization problems. Here, one wishes to find the ground state of a Hamiltonian $H_{P}$ that is diagonal in the computational basis and which encodes in its ground state the solution to some combinatorial problem of interest. The QAOA is constructed as $U(\bm{\theta})=\prod_{l}e^{-i\theta_{lM}H_{M}}e^{-i\theta_{lP}H_{P}}$, where $H_{M}$ is the so-called mixer Hamiltonian [119, 120, 121] usually taken to be $H=\sum_{j=1}^{n}X_{j}$. The presence of BPs in QAOA for maximum-cut tasks has been heuristically analyzed in [18, 122], where it has been shown that the loss will exhibit a BP for the vast majority of graphs if the circuit is deep enough. Moreover, this exponential concentration phenomenon has been also studied in QAOA variants such as the multi-angle QAOA where every gate in the circuit is assigned its own parameter [123]. Here, one can show that for all graphs with non-trivial cut value the circuit’s DLA dimension will be exponentially large implying that the loss will be concentrated [124]. Notably, the onset of BPs can even occur at depth-one for certain types of problems [124].
Unitary Coupled Singles and Doubles (UCCSD) ansatz. The UCCSD is a problem inspired PQC used to find the ground state of a fermionic molecular Hamiltonian via the variational quantum eigensolver [125, 13, 126]. Here, it has been shown that standard versions of UCC composed of controlled single excitation gates are particle-conserving universal unitaries [55]. Indeed, this implies that deep UCCSD ansatzes will exhibit a BP if the state of interest lies within a subspace with a number of excitations that scales with the problem size [127].

### v2

@@ III.3 Methods for analyzing variance
At the simplest level, one can determine if the loss or its gradients are concentrating by evaluating randomly sampled parameters and computing how the variance scales with the number of qubits [8, 31, 18]. Similarly, one can also compute other quantities proposed as surrogates to study whether a BP occurs These include the landscape’s information content [32] and fluctuation [33], the local state’s entanglement [34], or the landscape’s Fourier coefficients [35, 36, 37]. While these approaches can provide a rough estimate of scaling, they suffer from statistical errors due to finite measurements of the loss and generally cannot distinguish between the two types of BPs.
A preferred, but more theoretically demanding, method for detecting BPs is by analytically studying the scaling of the variance. In this method, assumptions about the PQC are necessary to make the problem more tractable. For instance, if $\mathcal{U}_{\bm{\theta}}$ is a deep unitary circuit whose distribution of unitaries forms at least a $2$-design [38] over a Lie group (for a brief introduction to the topic, see Box III.3 for Lie algebras, Box IV for an average over the group, and Box IV.2 for the PQC expressivity), one can use representation theory to exactly evaluate the variances [8, 39, 40, 41, 42, 18, 43, 44, 28, 45, 46, 47, 48] via Weingarten calculus [49, 50]. Instead, for shallow circuits that do not form designs, one can instead assume that the local gates are sampled from a local group, thus mapping the variance evaluation to a Markov chain-like process which can be analyzed analytically [51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63], numerically via Monte Carlo [64] and tensor networks [65, 66], or studied via XZ-calculus [67]. One can also take an iterative approach to explicitly integrate over the random parameters within a given angle range [68, 69, 59, 26] or exploit tools from Floquet theory [70, 71]. In addition, if $\mathcal{U}_{\bm{\theta}}$ is a noisy channel, one can study the variance by making assumptions regarding to the type of noise acting throughout the PQC [27, 72, 73, 74, 75].
@@ IV Origins of BPs
Theoretical and numerical calculations can help determine the flatness of the landscape, but they merely reveal symptoms of deeper issues in the variational quantum computing scheme. In this section we discuss recent findings which argue that BPs appear as a form of curse of dimensionality, as well as showcase how this phenomenon reveals itself under different scenarios. By understanding the connections between these different origins of BPs, we will see that one should be able to predict if a model will exhibit an exponentially concentrated loss.
@@ IV.1 A curse of dimensionality
Now, of course, this does not mean that comparing any pair of exponentially large objects will always lead to exponentially small signals. Indeed, many classical machine learning algorithms [81], as well as conventional quantum algorithms [82] manipulate vectors in the exponentially large Hilbert space by carefully choreographing the patterns of constructive and destructive interference of the probability amplitudes [83]. This is often because these methods incorporate specific structures or constraints that limit their expressiveness or control the exploration of the parameter space, thereby avoiding uncontrolled concentration effects. Moreover, it is also worth noting that conventional quantum algorithms have even been adapted to establish advantage proofs for quantum machine learning [84, 85]. Yet, there are not many such algorithms in the variational setting as their construction requires carefully and purposely designed circuits with appropriate inductive biases for the task they are trying to solve. How well variational quantum algorithms can be coerced into finding and exploiting such structure in the exponentially large Hilbert space is very much an open question.
@@ IV.2 Circuit expressiveness
[t] Box 6: Expressive power of the PQC The expressive power of a unitary PQC quantifies the breath of unitaries that the circuit is likely to produce in an unbiased manner. For instance, one can quantify the potential, or ultimate expressiveness of the PQC by its dynamical Lie algebra $\mathfrak{g}$ defined in Box III.3—since any unitary $U(\bm{\theta})$ produced by the PQC belongs in the dynamical Lie group $G=e^{\mathfrak{g}}$. Here, the larger the dimension of the algebra, the more expressive the circuit is, with controllable circuits being defined as those for which $\mathfrak{g}=\mathfrak{su}(2^{n})$. Of course, shallow circuits will not be able to cover the whole dynamical Lie group $G$, as they can only produce unitaries in some subset $\mathcal{S}\subseteq G$. To deal with this situation, we can instead consider the current expressiveness which is quantified by the difference between $\tau_{\mathcal{S}}^{(t)}$ and $\tau_{G}^{(t)}$ (where we recall that these quantities are defined in Box IV). When those operators match, we say that the PQC forms a $t$-design [38]. Importantly, if $|\tau_{\mathcal{S}}^{(2)}-\tau_{G}^{(2)}|\leqslant\varepsilon$ (in operator norm) then [43] $|{\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]-{\rm Var}_{G}[\ell_{\bm{\theta}}(\rho,O)]|\in\mathcal{O}(\varepsilon\lVert O\rVert_{1})\,,$ (12) showing that we can quantify how much the variance of a shallow circuit deviates from that of a Haar random PQC over $G$ via $\varepsilon$.
Perhaps the main culprit responsible for the curse of dimensionality is the PQC as it has been shown that its expressive power (see Box IV.2) is directly linked to BPs. Roughly speaking, the larger the space the PQC explores in an unbiased manner, the more it is prone –when randomly initialized– to lead to input states that have exponentially small inner products with the measurement operator. The previous connection between expressiveness and BPs has been mathematically formalized for unitary [8, 51, 86, 18, 43, 44, 28, 87] and noisy [88] circuits, as one can show that the more expressive a circuit is, the more the loss can concentrate. Similarly, the PQC’s capacity  [89, 90] –measured by how many functions it can represent–affects the concentration of the loss. The higher the capacity, the more likely it is that the loss will concentrate [91].
@@
In particular, let us review the recent approach in Refs. [43, 44, 28] for obtaining an exact analytical form for the loss function variance of deep unitary PQCs. By deep we refer to the regime where the number of layers is large enough to guarantee that the distribution of unitaries corresponding to random parameter choices forms a $2$-design [38]. That is, that it approximately matches the Haar distribution on $e^{\mathfrak{g}}$ (where $\mathfrak{g}$ is the circuit’s dynamical Lie algebra (DLA), defined in Box III.3) up to second moments with small additive approximation error. Assume for simplicity that either $\rho$ or $O$ belong to a group-module $\mathcal{M}\subseteq\mathcal{B}(\mathcal{H})$ that is non-trivial under the adjoint action of the circuit. (We recall that a group-module is a vector space closed under the action of the group. In the case of $\mathcal{B}$, $\mathcal{M}$ is a subspace that is closed under the adjoint action of $e^{\mathfrak{g}}$, i.e., if $A\in\mathcal{M}$, then $UAU^{\dagger}$ also belongs to $\mathcal{M}$ for any $U\in e^{\mathfrak{g}}$  [^1] Technically speaking, group modules are irreducible representations of the dynamical Lie group over the vector space $\mathcal{B}(\mathcal{H}_{0})$). One finds
@@
Here, $P_{\mathcal{M}}(\rho)$ and $P_{\mathcal{M}}(O)$ are the norms of the projection of $\rho$ and $O$ onto the module $\mathcal{M}$. Explicitly, given an Hermitian orthonormal basis for $\mathcal{M}$, denote as $\{H_{\mu}\}_{\mu}$, then $P_{\mathcal{M}}(\rho)=\sum_{\mu}{\rm Tr}[H_{\mu}\rho]^{2}$. As one can see, and as further discussed below, these three quantities on the right hand side of Eq. (13) are associated with different in which BPs can arise. In particular, expressive power of PQC is related to $\dim({\mathcal{M}})$ and the role of an input state as well as measurements are associated with $P_{\mathcal{M}}(\rho)$ and $P_{\mathcal{M}}(O)$, respectively. Having a PQC with exponentially large $\dim({\mathcal{M}})$ and/or exponentially small $P_{\mathcal{M}}(\rho)$ and/or $P_{\mathcal{M}}(O)$ results in an exponentially small variance and hence BPs.
@@ IV.3 Input states and measurements
While highly expressive PQCs significantly contribute to BPs, circuits with limited expressiveness can also exhibit BPs depending on the initial state and measurement operator. For instance, let us consider Eq. (13) again, and assume that the ensuing DLA is not exponentially large, i.e., that we have $\dim(\mathfrak{g})\in\mathcal{O}(\operatorname{poly}(n))$. This occurs when $\mathcal{U}_{\bm{\theta}}$ has strong inductive biases such as encoding some symmetries (see below for specific examples). Now, despite using a circuit with a non-exponentially large underlying algebra, the loss can still be exponentially concentrated. For instance, if we pick $O$ belonging to an exponentially large module (i.e., small algebras can admit exponentially large modules), then one will find exponential concentration from the denominator in Eq. (13). However, if $O$ belongs to a module whose dimension grows polynomially with $n$, such as $iO\in\mathfrak{g}$, the variance’s denominator $\dim(\mathcal{M})=\dim(\mathfrak{g})$ will only decrease polynomially with $n$, meaning that the circuit expressiveness cannot be responsible for a BP. In this case, the loss is obtained by comparing objects in a polynomially large subspace, potentially avoiding the curse of dimensionality. However, a deterministic BP is still possible if the norm of the projection of $\rho$ into the module is exponentially small, i.e., if $\mathcal{P}_{\mathcal{M}}(\rho)\in\mathcal{O}(1/2^{n})$. Box IV.3 presents an example where an inexpressive circuit can still lead to a BP, either because $O$ belongs to an exponentially large $\mathcal{M}$ or because $O$ comes from a polynomial module but the state is misaligned.
@@
[t] Box 7: Single-qubit rotation PQCs can have BPs Consider a unitary circuit composed of general single qubit gates as $U(\bm{\theta})=\prod_{j=1}^{n}U_{j}(\bm{\theta}_{j})$, where $U_{j}(\bm{\theta}_{j})$ acts only on the $j$-th qubit. Clearly, this circuit is fairly inexpressive as it does not even generate entanglement. Moreover, its DLA is simply given by $\mathfrak{g}=\oplus_{j=1}^{n}\mathfrak{su}(2)$, and thus of dimension $\dim(\mathfrak{g})=3n\in\mathcal{O}(\operatorname{poly}(n))$. Yet, we can see that if $O=Z^{\otimes n}$, then the operator belongs to an exponentially large module composed of all $3^{n}$ Pauli operators acting non trivially on all the qubits. Assuming that each $U_{j}(\bm{\theta}_{j})$ forms a $2$-design over $\mathbb{SU}(2)$ [43], a direct calculation reveals that for all $\rho$, ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]\in\mathcal{O}(1/2^{n})$. However, if $O=Z_{1}$, then the measurement belongs to a module composed of the $3$ Pauli operators acting on the first qubit. Now, Eq. (13) leads to ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]=2({\rm Tr}[\rho_{1}^{2}]-1/2)/3$, with $\rho_{1}$ the reduced state of $\rho$ on the first qubit. The choice of measurement operator can lead to a BP. When $O=Z^{\otimes n}$ the loss is obtained by comparing operators that can move in exponentially large modules, while for $O=Z_{1}$, the circuit can only move in the subspace of operators acting only on the first qubit. However, while the PQC and measurements are not responsible for BPs, the initial state can lead to an exponentially concentrated loss. In particular, if the state is too entangled and $({\rm Tr}[\rho_{1}^{2}]-1/2)\in\mathcal{O}(1/2^{n})$ (which occurs if $\rho$ satisfies a volume law of entanglement [29]), then a deterministic BP occurs.
@@ V Barren-plateau-prone architectures
In the previous section we have provided some intuition regarding the origins of BPs. In this section we review some BP-prone architectures. We note that in most cases, deep versions of these architectures can lead to exponentially concentrated losses due to their large DLAs. However, these architectures can still be useful in shallow circuits or when using smart initialization strategies.
@@ V.1 Hardware efficient ansatz and deep unstructured circuits
Hardware efficient ansatz [116] is a generic term employed for unstructured PQCs composed of single-qubit rotation gates interleaved with (fixed or parametrized) entangling gates. The term hardware-efficient spans from the fact that the gates of the circuit are chosen not because they encode some strong inductive bias from the problem one is trying to solve, but because they are easy to implement on the device (e.g., one only picks entangling gates that follow the device connectivity to avoid compiling overhead [95, 117, 118]). Under quite general conditions, hardware efficient ansatzes will be universal  [18], meaning that their DLA is $\mathfrak{g}=\mathfrak{su}(2^{n})$. As such, deep versions of these circuits are as expressive as possible, and therefore will exhibit BPs irrespective of the initial state or measurement operator [119, 39, 8, 51, 120, 121, 122]. Here it is important to note that several standard architectures exist in the literature, such as entangling gates acting on a brick-like fashion on alternating pairs of qubits or on a ladder. While these can indeed exhibit different performances in their shallower forms, they will still lead to a BP if the number of layers becomes too large.
@@ V.2 Some problem-inspired ansatzes
Just like hardware efficient is used for a wide-range of problem-agnostic circuits, problem-inspired ansatzes is a blanket term used for models which embed inductive biases about the problem at hand into the circuit architecture itself. In most cases, problem-inspired ansatzes adjust circuit expressiveness by using specific sets of gates, distinct entangling gate topologies, or by aligning parameters with the symmetries of the task. Here, one can generally expect that $\mathfrak{g}$ will be smaller than $\mathfrak{su}(2^{n})$ [43]. While we will present below circuits which can in fact sufficiently reduce the circuit expressive power to potentially avoid BPs, we here discuss problem-inspired architectures that can still have exponentially concentrated loss functions despite their inductive biases.
@@
Quantum Alternating Operator Ansatz (QAOA). The QAOA [125, 126] is widely used to variationally attempt to solve combinatorial optimization problems. Here, one wishes to find the ground state of a Hamiltonian $H_{P}$ that is diagonal in the computational basis and which encodes in its ground state the solution to some combinatorial problem of interest. The QAOA is constructed as $U(\bm{\theta})=\prod_{l}e^{-i\theta_{lM}H_{M}}e^{-i\theta_{lP}H_{P}}$, where $H_{M}$ is the so-called mixer Hamiltonian [127, 128, 129] usually taken to be $H=\sum_{j=1}^{n}X_{j}$. The presence of BPs in QAOA for maximum-cut tasks has been analyzed in [18, 130, 131], where it has been shown that the loss will exhibit a BP for the vast majority of graphs if the circuit is deep enough. Nevertheless, it is an open question the depth for which QAOA can obtain useful solution and outperform state-of-the-art classical methods [132, 22, 133, 134]. Moreover, this exponential concentration phenomenon has been also studied in QAOA variants such as the multi-angle QAOA where every gate in the circuit is assigned its own parameter [135]. Here, one can show that for all graphs with non-trivial cut value the circuit’s DLA dimension will be exponentially large implying that the loss will be concentrated [131, 136, 137, 138]. Notably, while the depth at which BPs will occur depends on the graph, and will likely require a polynomial number [43, 139], there exists cases where QAOA can have BPs even with a depth of one [131]. With the previous being said, there have been promising research direction to train QAOA in regimes where the BP proofs do not hold (e.g., training classically on small problems and transferring parameters)  [140, 141, 142, 143, 144, 145]. However, the performance of these approaches at scale is still an open question.
Unitary Coupled Singles and Doubles (UCCSD) ansatz. The UCCSD is a problem inspired PQC used to find the ground state of a fermionic molecular Hamiltonian via the variational quantum eigensolver [146, 13, 147]. Here, it has been shown that standard versions of UCC composed of controlled single excitation gates are particle-conserving universal unitaries [60]. Indeed, the results in Ref. [60] implies that deep randomly initialized UCCSD ansatzes will exhibit a BP if the state of interest lies within a subspace with a number of excitations that scales with the problem size (see [148] for a numerical verification of this fact). In the next section we discuss how some of these issues could be avoided, e.g., by using shallow circuits, and good initializations strategies.